OpenStudy (anonymous):

7 females, and 5 males including Larry. There are 4 tasks to be assigned. Note that assigning the same people different tasks constitutes a different assignment. (1) Find the probability that both males and females are given a task. (2) Find the probability that Larry and at least one female are given tasks.

Directrix (directrix):

Hey, CW.

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

wow check this out http://www.jiskha.com/display.cgi?id=1236051521

OpenStudy (anonymous):

larry must be a popular guy

OpenStudy (anonymous):

how ya doin g

OpenStudy (anonymous):

doing

OpenStudy (anonymous):

these word problems are killing me

OpenStudy (anonymous):

it is the same problem with different numbers

OpenStudy (anonymous):

I had to take an exam today. And this homework set is due on tomorrow at 730 am. :)

OpenStudy (anonymous):

2 days to complete them

OpenStudy (anonymous):

no i am not....

OpenStudy (anonymous):

I like to know what i am doing though

Directrix (directrix):

Wyatt Earp: Speed is great but accuracy is everything.

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

i am always mixing these up as I work through them

OpenStudy (anonymous):

I have 12 or 13 problems due in the morning

OpenStudy (anonymous):

I feel like a math bum :)

Directrix (directrix):

Don't feel. Just work. Post the other 12 or 13 problems individually on this site. You do that while I work on this problem. I may have to stop and eat some brain food. :} You need to focus. Nobody here is a bum.

OpenStudy (anonymous):

is that the final solution

OpenStudy (anonymous):

for both male and female

OpenStudy (anonymous):

i know I am very lost haha

OpenStudy (anonymous):

it didnt work though

OpenStudy (anonymous):

why the zero in both combinations

OpenStudy (anonymous):

not quite brain food :) and this one

OpenStudy (anonymous):

ha ha i post them wuickly

OpenStudy (anonymous):

quickly

Directrix (directrix):

P(both genders) = 1 - { [ C(7,0) C(5,4) - C(7,4) C(5,0)] / C(12,4) } = 1 - 40/495 = 455/495 = 91/99 = .919 approx

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

So the second part would be similar?

Directrix (directrix):

P(Larry and at least 1 F) P(at least 1F) = 1 - P(0F) = 1 -{ [C(7,0)C(5,4)] } / C(12,4) = 5/495 = 490/495 = 98/99 = .99 ish. Ways to choose Larry = 1 C(1,1) = 1 P(Larry and at least 1 F) = 1(98/99) = 98/99 = .99 approx

OpenStudy (anonymous):

no it didnt work

OpenStudy (anonymous):

larry wouldnt be 5?

OpenStudy (anonymous):

5,1

Directrix (directrix):

But, we want Larry and none of the other males. So P(Larry) = C(1,1) C(4,0) = 1. That would be Larry and none of the other males.

Directrix (directrix):

I'm going to cases for this. Slow but may make more sense.

OpenStudy (anonymous):

how ya doing

Directrix (directrix):

Larry and at least 1F -------- Cases {Lar, 1F, 2 other M}; {Lar, 2F, 1 other M}; and {Lar, 3F, 0 other M}

Directrix (directrix):

Ways to choose Lar, 1F, 2 other M} = 1 C(7,1) C(4,2) = 42 Ways to choose {Lar, 2F, 1 other M}; = 1 C(7,2) C(4,1)= 84 Ways to choose {Lar, 3F, 0 other M} = 1 C(7,3) C(4,0) = 35 Ways to choose 4 from 12 is C(12,4) = 495 P(L and at least 1F) = [42 + 84 +35] / 495 = 161/495 =.325 approx

OpenStudy (anonymous):

CORRECT! You are a genius.... I have posted some more. I haven't wanted to post to many at a time and flood the screen

OpenStudy (anonymous):

I am wrong no matter what i do on these

Directrix (directrix):

Let's find another problem.

OpenStudy (anonymous):

I used a lot of the work i learned from hear that helped me on the test in some of these.